# MI-NeRF: Learning a Single Face NeRF from Multiple Identities

### v1

@@ 4.1 Dataset and Evaluation
Dataset. To evaluate our proposed method, we collected 140 talking face videos of different identities from publicly available datasets [23, 35, 8, 24, 19, 1, 14, 69, 63]. We included a variety of standard front-facing videos (e.g. political speeches), as well as more challenging videos, with large variations in head pose, lighting, and expressiveness (e.g. movies and news satire television programs). The videos are in HD quality (720p) and of around 20 seconds to 5 minutes duration. For each video, we run the 3DMM fitting procedure, as described in Sec. 3.1. We use 100 videos for our training set and we keep the rest as novel identities. We keep the last 10% of the frames of the training videos as our test set. More details of our dataset collection are included in the appendix.
Evaluation Metrics. We measure the visual quality of the generated videos, using peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [64], and learned perceptual image patch similarity (LPIPS) [68], and we verify the identity of the target subject, using the average content distance (ACD) [57, 53]. Additionally, we use the LSE-D (Lip Sync Error - Distance) and LSE-C (Lip Sync Error - Confidence) metrics [43, 12], to assess the lip synchronization, i.e., if the generated expressions are meaningful given the corresponding speech signal.
@@ 4.2 Ablation Study
We conduct an ablation study on the multiplicative module and the input conditions of MI-NeRF. Given 10 videos from different identities, we investigate variants of our model (see Tab. 1). After training, we synthesize a video of each identity given input expressions from another one. In this way, we evaluate if the model learns to disentangle identity and non-identity specific information.
Firstly, we evaluate the possibility of simply concatenating all the input vectors, as usually done in standard identity-specific NeRFs [15, 23, 8], i.e., omitting the multiplicative module. We call this “Baseline NeRF”. As shown in Tab. 1 and Fig. 3, Baseline NeRF cannot learn to disentangle between different identities. Based on the identity-expression pairs seen during training, it frequently synthesizes a different identity than the target one, or a mixture of identities, leading to visible artifacts.
Variants of the proposed multiplicative module $M$ are examined in rows (A1) to (A7) to determine the simplest and most effective option. We started by just learning a simple mapping of the $\bm{e}$ and $\bm{i}$ vectors, using $\bm{W}_{2}$ and $\bm{W}_{3}$ matrices (A1). We then explored variants that include a Hadamard product $(\bm{e}*\bm{i})$ to model their non-linear interaction. The last variant (A7) is inspired by Wang et al. [61] and uses $\bm{\mathcal{W}}\in\mathbb{R}^{d\times d\times d}$, requiring significantly more parameters than our proposed $M$ ($d^{3}$ vs $d^{2}$). The variant (A6) can be derived from (A7) (see appendix). All these variants (A1) - (A7) lead to different disentanglement results, with many of them being quite poor.
@@
Figure 3: Ablation Study. Qualitative comparison of MI-NeRF with Baseline NeRF that concatenates all input conditions, without using any multiplicative module, that leads to poor disentanglement. Our proposed multiplicative module demonstrates robustness, disentangling between identity and expression.
@@ 4.3 Facial Expression Transfer
In this section, we demonstrate the effectiveness of MI-NeRF for facial expression transfer. Given an identity from the training set, MI-NeRF enables explicit control of their expressions, and thus can synthesize high-quality videos of them given novel expressions as input (see Fig. 4). It can animate multiple identities under the same novel expression and novel view (see Fig. 5).
Comparisons. Fig. 6 demonstrates the results of MI-NeRF with $M$ (personalized) for challenging novel expressions, not seen for the target identity. We compare with methods that similarly allow meaningful control of expression parameters. NeRFace [15] is a standard identity-specific NeRF, without any multiplicative module. Its immediate extension is the Baseline NeRF that is trained on multiple identities, concatenating additional identity vectors as input. INSTA [72] learns a single-identity NeRF, embedded around FLAME [31] and based on neural graphics primitives, that enables faster training and rendering. HeadNeRF [25] is a NeRF-based parametric head model, trained on a large amount of high-quality images from multiple identities. Tab. 2 shows the corresponding quantitative results of facial expression transfer, computed on 10 synthesized videos from our test set.
We notice that NeRFace and INSTA can produce good visual quality, but cannot generalize to unseen expressions. Trained only on a video of a single identity, they have only seen a limited variety of facial expressions, and thus can lead to artifacts in case of novel input expressions (see Fig. 6). HeadNeRF distorts the target identity and inaccurately produces the target facial expression. Our method demonstrates robustness in synthesizing novel expressions for any identity, as it leverages information from multiple subjects during training. Fig. 7(b) shows the corresponding PSNR, computed on frames with challenging unseen expressions for a particular subject. MI-NeRF significantly outperforms the standard single-identity NeRFace [15] and INSTA [72] in such cases. Increasing the number of training identities improves the robustness of our generic model. Further personalization enhances the final visual quality.
Training Time. MI-NeRF significantly outperforms the standard single-identity NeRF-based methods, like NeRFace [15], in terms of the training time (see Fig. 7(a)). Training on 100 identities needs only about 80 hours of training, compared to 40 hours per identity needed by standard NeRFs, leading to a $90\%$ decrease approximately. Further personalization for a target identity adds only another 5-8 hours of training on average, enhancing the visual quality (see Fig. 7(b)). In the appendix, we show that our multiplicative module can also be applied to faster methods, like INSTA [72], extending them to multiple identities and similarly achieving a $90\%$ decrease in training time for 100 identities. We encourage the readers to watch the supplementary video.
Figure 6: Transferring Novel Expressions. Qualitative comparison of MI-NeRF with state-of-the-art approaches when transferring unseen expressions to a target identity. NeRFace [15] is a single-identity NeRF, INSTA [72] is a single-identity geometry-guided deformable NeRF, and HeadNeRF [25] is a NeRF-based parametric head model trained on a large dataset. Our method demonstrates robustness in synthesizing novel (unseen) expressions for any input identity.
(a) Total training time vs number of identities
(b) Visual quality vs number of identities
Figure 7: (a) Total training time vs total number of identities. Standard single-identity NeRFs, like NeRFace [15], AD-NeRF [23], and LipNeRF [8] require approximately 40 hours training per identity. On the contrary, our MI-NeRF (generic) can be trained on 100 identities in 80 hours, leading to a $90\%$ decrease approximately. Further personalization takes another 5-8 hours per identity. (b) Corresponding visual quality of generated videos with challenging novel expressions, measured by PSNR (higher the better). Increasing the number of identities improves the robustness of our model to unseen expressions.
Figure 8: Lip Synced Video Synthesis. Qualitative comparison of our method with state-of-the-art approaches, GAN-based Wav2Lip [43], AD-NeRF [23] and Lip-NeRF [8]. The original video is in English (1st column). The generated videos (columns 2-5) are lip synced to dubbed audio in Spanish.
@@
NeRFace [15] 28.11 0.88 0.13 0.11
INSTA [72] 28.02 0.87 0.12 0.10
HeadNeRF [25] 19.89 0.74 0.32 0.85
@@
AD-NeRF [23] 11.40 1.28 27.38 0.84
LipNeRF [8] 9.92 2.71 30.10 0.89
GeneFace [66] 11.80 2.22 25.56 0.85
@@ 4.4 Lip Synced Video Synthesis
MI-NeRF can be also used for audio-driven talking face video synthesis, also called lip-syncing. Prior work, LipNeRF [8], has shown that conditioning a NeRF on the 3DMM expression space, compared to audio features, leads to more accurate and photorealistic lip synced videos. Our method naturally extends LipNeRF to multiple identities, using their proposed audio-to-expression mapping. For this task, we used the dataset proposed by LipNeRF [8]  [^2] We would like to thank the authors of LipNeRF for providing the cinematic data from YouTube.. It includes 10 videos from popular movies in English, of around 30 seconds to 2 minutes long each in HD resolution (720p), and corresponding dubbed audio in 2 or 3 different languages for each video, including French, Spanish, German and Italian.
Results. Tab. 3 shows the corresponding quantitative evaluation of the synthesized videos, lip synced to dubbed audio in different languages. Leveraging information from multiple identities, MI-NeRF slightly improves the LSE metrics and achieves similar visual quality with LipNeRF, at much lower training time (see Fig. 7). AD-NeRF [23] and GeneFace [66] lack in lip synchronization for this task. Fig. 8 compares the mouth position for two examples, lip synced to Spanish (original audio in English). The GAN-based method, Wav2Lip [43], frequently produces artifacts and blurry results. Since it operates in the 2D image space, it cannot handle large 3D movements. AD-NeRF overfits to the training audio and cannot generalize well to different speech inputs. LipNeRF performs well, but sometimes cannot produce unseen expressions (e.g. does not close the mouth in the first row), since it is only trained on a single video of limited duration. Both AD-NeRF and LipNeRF are standard single-identity NeRFs, requiring expensive identity-specific optimization.
@@ 4.5 Short-Video Personalization
MI-NeRF can also be adapted to an unseen identity, i.e., that is not seen as a part of the initial training set (see Sec. 3.4). Fig. 15 demonstrates the results when only a very short video of the new identity is available (1 or 3 seconds length). Please note that in this case we use a small number of consecutive frames, and thus a very small part of the expression space is covered. In this case, the single-identity INSTA [72] cannot learn a good head representation of the identity, given that very short video, and produces artifacts. HeadNeRF is trained on a large dataset of images and its fitting requires only a single image. However, it fails to capture the identity accurately. Our method demonstrates robustness under unseen expressions and novel views for the new identity. We include more quantitative and qualitative results for our short-video personalization in the supplementary. We also encourage the readers to watch our supplementary video.
Figure 9: Short-Video Personalization. Learning a novel identity from a short video of 1 or 3 seconds duration. Qualitative comparison of MI-NeRF with the single-identity INSTA [72] and the NeRF-based parametric head model HeadNeRF [25]. Our method demonstrates robustness across unseen expressions and novel views for any new identity.
@@ 5 Conclusion
In this work, we introduce MI-NeRF that learns a single dynamic NeRF from monocular talking face videos of multiple identities. We propose a multiplicative module that captures the non-linear interactions of identity and non-identity specific information. Trained on multiple identities, MI-NeRF significantly reduces the training time over standard single-identity NeRFs. Our model can be further personalized for a target identity, given only a short video of a few seconds length, achieving state-of-the-art performance for facial expression transfer and talking face video synthesis. In the future, we envision extending our approach to thousands of identities, learning collectively from very short in-the-wild videos.
Appendix
The appendix is organized as follows:
  • Technical Derivations of the Modules in Sec. A.
  • Additional Ablation Study in Sec. B.
  • Additional Results in Sec. C.
  • Comparison with Faster Methods in Sec. D.
  • Implementation Details in Sec. E.
  • Limitations in Sec. F.
  • Ethical Considerations in Sec. G.
  • Dataset Details in Sec. H.
We strongly encourage the readers to watch our supplementary video.
A Technical analysis of the models
@@
Notation. Tensors are symbolized by calligraphic letters, e.g., $\bm{\mathcal{X}}$. The mode-$m$ vector product of $\bm{\mathcal{X}}$ with a vector $\bm{u}\in\mathbb{R}^{I_{m}}$ is denoted by $\bm{\mathcal{X}}\times_{m}\bm{u}$. A core tool in our analysis is the CP decomposition [28]. By considering the mode-$1$ unfolding of an $M^{\text{th}}$-order tensor $\bm{\mathcal{X}}$, the CP decomposition can be written in matrix form as in [28]:
@@
A.1 Proof of Proposition 1
To prove the Proposition 1 (see Sec. 3.2 of the main paper), we will construct the special form from the general multiplicative interaction.
@@ A.1 Proof of Proposition 1
where $\bm{W}_{2},\bm{W}_{3}\in\mathbb{R}^{o\times d}$ are learnable parameters. We can use the CP decomposition [28] to induce a low-rank decomposition on the third-order tensor $\bm{\mathcal{W}}$. This results in the following expression:
@@
where * is the Hadamard product. If we set $o=k=d$ and set $A^{T}=\bm{U}_{1},B^{T}=\bm{U}_{2}$, then we end up with the model of Eq. (1) of the main paper, which concludes the proof.
A.2 Intuition on Equation (2)
The model of Eq. (2) can also capture multiplicative interactions in case of $N=2$ or high-degree interactions in case $N>2$. We provide some constructive proof to show the case of $N=2$ below, since the number of terms increases fast already for this case. Subsequently, we focus on the case of high-degree interactions.
Proposition 2
For $N=2$, the model of Eq. (2) captures multiplicative interactions between the expression $\bm{e}$ and the identity $\bm{i}$ vector.
Proof
For $N=2$, Eq. (2) becomes:
$\begin{split}H(\bm{e},\bm{i})=\\ \bm{C}\left\{\left(\bm{U}_{(2,1)}\bm{e}+\bm{U}_{(2,2)}\bm{i}\right)*\left(\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}\right)\right\}+\bm{C}\left(\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}\right)=\\ \bm{C}\left\{(\bm{U}_{(2,1)}\bm{e})*(\bm{U}_{(1,1)}\bm{e})\right\}+\bm{C}\left\{(\bm{U}_{(2,1)}\bm{e})*(\bm{U}_{(1,2)}\bm{i})\right\}+\\ \bm{C}\left\{(\bm{U}_{(2,2)}\bm{i})*(\bm{U}_{(1,1)}\bm{e})\right\}+\bm{C}\left\{(\bm{U}_{(2,2)}\bm{i})*(\bm{U}_{(1,2)}\bm{i})\right\}+\\ \bm{C}\bm{U}_{(1,1)}\bm{e}+\bm{C}\bm{U}_{(1,2)}\bm{i}\;.\end{split}$ (11)
Each of the first terms of the last equation arises from a CP decomposition with specific factor matrices (the inverse process from Eq. 8 to Eq. 9 can be followed). Therefore, Eq. 11 captures multiplicative interactions between the two variables, including second-degree interactions among the same variable.
Proposition 3
For $N>2$, the model of Eq. (2) captures high-degree interactions between the expression $\bm{e}$ and the identity $\bm{i}$ vector.
Proof
The number of terms increase rapidly in Eq. (2). Without loss of generality, (a) we will only use the multiplicative term (and ignore the additive term of $+\bm{x}_{n-1}$, since this is not contributing to the higher-degree terms) and (b) we will showcase this for $N=3$. For $N=3$, we obtain the following formula:
@@ A.2 Intuition on Equation (2)
Following similar arguments as the proofs above and the unfolding of the CP decomposition, we can exhibit that all of those terms arise from CP decompositions with specific factor matrices, while each one captures a triplet of the form $(\bm{\tau}_{1},\bm{\tau}_{2},\bm{\tau}_{3})$ with $\bm{\tau}\in\{\bm{e},\bm{i}\}$.
B Additional Ablation Study

### v2

@@ 4.1 Dataset and Evaluation
Dataset. To evaluate our proposed method, we collected 140 talking face videos of different identities from publicly available datasets (Guo et al., 2021; Lu et al., 2021; Chatziagapi et al., 2023; Hazirbas et al., 2021; Ginosar et al., 2019; Ahuja et al., 2020; Duarte et al., 2021; Zhang et al., 2021; Wang et al., 2021b). We included a variety of standard front-facing videos (e.g. political speeches), as well as more challenging videos, with large variations in head pose, lighting, and expressiveness (e.g. movies and news satire television programs). The videos are in HD quality (720p) and of around 20 seconds to 5 minutes duration. For each video, we run the 3DMM fitting procedure, as described in Sec. 3.1. We use 100 videos for our training set and we keep the rest as novel identities. We keep the last 10% of the frames of the training videos as our test set. More details of our dataset collection are included in the appendix.
Evaluation Metrics. We measure the visual quality of the generated videos, using peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) (Wang et al., 2004), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), and we verify the identity of the target subject, using the average content distance (ACD) (Vougioukas et al., 2019; Tulyakov et al., 2018). Additionally, we use the LSE-D (Lip Sync Error - Distance) and LSE-C (Lip Sync Error - Confidence) metrics (Prajwal et al., 2020; Chung & Zisserman, 2016), to assess the lip synchronization, i.e. if the generated expressions are meaningful given the corresponding speech signal.
@@
Figure 2: Ablation Study. Qualitative comparison of MI-NeRF with Baseline NeRF that concatenates all input conditions, without using any multiplicative module, and leads to poor disentanglement. Our proposed multiplicative module demonstrates robustness, disentangling between identity and expression.
Figure 3: Transferring Novel Expressions. Qualitative comparison of MI-NeRF with state-of-the-art approaches when transferring unseen expressions to a target identity. NeRFace (Gafni et al., 2020) is a single-identity NeRF, INSTA (Zielonka et al., 2023) is a single-identity geometry-guided deformable NeRF, and HeadNeRF (Hong et al., 2022) is a NeRF-based parametric head model trained on a large dataset. Our method demonstrates robustness in synthesizing novel (unseen) expressions for any input identity.
@@ 4.2 Ablation Study
We conduct an ablation study on the multiplicative module and the input conditions of MI-NeRF. Given 10 videos from different identities, we investigate variants of our model (see Table 1). After training, we synthesize a video of each identity given input expressions from another one. In this way, we evaluate if the model learns to disentangle identity and non-identity specific information.
Firstly, we evaluate the possibility of simply concatenating all the input vectors, as usually done in standard identity-specific NeRFs (Gafni et al., 2020; Guo et al., 2021; Chatziagapi et al., 2023), i.e. omitting the multiplicative module. We call this “Baseline NeRF”. As shown in Table 1 and Fig. 2, Baseline NeRF cannot learn to disentangle between different identities. Based on the identity-expression pairs seen during training, it frequently synthesizes a different identity than the target one, or a mixture of identities, leading to visible artifacts.
Variants of the proposed multiplicative module $M$ are examined in rows (A1) to (A7) to determine the simplest and most effective option. We started by just learning a simple mapping of the $\bm{e}$ and $\bm{i}$ vectors, using $\bm{W}_{2}$ and $\bm{W}_{3}$ matrices (A1). We then explored variants that include a Hadamard product $(\bm{e}*\bm{i})$ to model their non-linear interaction. The last variant (A7) is inspired by Wang et al. (2019) and uses $\bm{\mathcal{W}}\in\mathbb{R}^{d\times d\times d}$, requiring significantly more parameters than our proposed $M$ ($d^{3}$ vs $d^{2}$). The variant (A6) can be derived from (A7), using similar arguments to Sec. A.1. All these variants (A1) - (A7) lead to different disentanglement results, with many of them being quite poor.
@@ 4.3 Facial Expression Transfer
In this section, we demonstrate the effectiveness of MI-NeRF for facial expression transfer. Given an identity from the training set, MI-NeRF enables explicit control of their expressions, and thus can synthesize high-quality videos of them given novel expressions as input.
Figure 4: Left: Total training time vs total number of identities. Standard single-identity NeRFs, like NeRFace (Gafni et al., 2020), AD-NeRF (Guo et al., 2021), and LipNeRF (Chatziagapi et al., 2023) require approximately 40 hours training per identity. On the contrary, our MI-NeRF (generic) can be trained on 100 identities in 80 hours, leading to a $90\%$ decrease approximately. Further personalization takes another 5-8 hours per identity. Right: Corresponding visual quality of generated videos with challenging novel expressions, measured by PSNR (higher the better). Increasing the number of identities improves the robustness of our model to unseen expressions.
@@
NeRFace (Gafni et al., 2020) 28.11 0.88 0.13 0.11
INSTA (Zielonka et al., 2023) 28.02 0.87 0.12 0.10
HeadNeRF (Hong et al., 2022) 19.89 0.74 0.32 0.85
@@
Comparisons. Fig. 3 demonstrates the results of MI-NeRF with $M$ (personalized) for challenging novel expressions, not seen for the target identity. We compare with methods that similarly allow meaningful control of expression parameters. NeRFace (Gafni et al., 2020) is a standard identity-specific NeRF, without any multiplicative module. Its immediate extension is the Baseline NeRF that is trained on multiple identities, concatenating additional identity vectors as input. INSTA (Zielonka et al., 2023) learns a single-identity NeRF, embedded around FLAME (Li et al., 2017) and based on neural graphics primitives, that enables faster training and rendering. HeadNeRF (Hong et al., 2022) is a NeRF-based parametric head model, trained on a large amount of high-quality images from multiple identities. Table 2 shows the corresponding quantitative results of facial expression transfer, computed on 10 synthesized videos from our test set.
We notice that NeRFace and INSTA can produce good visual quality, but cannot generalize to unseen expressions. Trained only on a video of a single identity, they have only seen a limited variety of facial expressions, and thus can lead to artifacts in case of novel input expressions (see Fig. 3). HeadNeRF distorts the target identity and inaccurately produces the target facial expression. Our method demonstrates robustness in synthesizing novel expressions for any identity, as it leverages information from multiple subjects during training. Fig. 4 (right) shows the corresponding PSNR, computed on frames with challenging unseen expressions for a particular subject. MI-NeRF significantly outperforms the standard single-identity NeRFace (Gafni et al., 2020) and INSTA (Zielonka et al., 2023) in such cases. Increasing the number of training identities improves the robustness of our generic model. Further personalization enhances the final visual quality.
Training Time. MI-NeRF significantly outperforms the standard single-identity NeRF-based methods, like NeRFace (Gafni et al., 2020), in terms of the training time (see Fig. 4 left). Training on 100 identities needs only about 80 hours of training, compared to 40 hours per identity needed by standard NeRFs, leading to a $90\%$ decrease approximately. Further personalization for a target identity adds only another 5-8 hours of training on average, enhancing the visual quality (see Fig. 4 right). In the appendix, we show that our multiplicative module can also be applied to faster methods, like INSTA (Zielonka et al., 2023), extending them to multiple identities and similarly achieving a $90\%$ decrease in training time for 100 identities. We encourage the readers to watch our suppl. video.
Figure 5: Lip Synced Video Synthesis. Qualitative comparison of our method with state-of-the-art approaches, GAN-based Wav2Lip (Prajwal et al., 2020), AD-NeRF (Guo et al., 2021) and Lip-NeRF (Chatziagapi et al., 2023). The original video is in English (1st column). The generated videos (columns 2-5) are lip synced to dubbed audio in Spanish.
@@
AD-NeRF (Guo et al., 2021) 11.40 1.28 27.38 0.84
LipNeRF (Chatziagapi et al., 2023) 9.92 2.71 30.10 0.89
GeneFace (Ye et al., 2023) 11.80 2.22 25.56 0.85
@@ 4.4 Lip Synced Video Synthesis
MI-NeRF can be also used for audio-driven talking face video synthesis, also called lip-syncing. Prior work, LipNeRF (Chatziagapi et al., 2023), has shown that conditioning a NeRF on the 3DMM expression space, compared to audio features, leads to more accurate and photorealistic lip synced videos. Our method naturally extends LipNeRF to multiple identities, using their proposed audio-to-expression mapping. For this task, we used the dataset proposed by LipNeRF (Chatziagapi et al., 2023)  [^2] We would like to thank the authors of LipNeRF for providing the cinematic data from YouTube.. It includes 10 videos from popular movies in English, of around 30 seconds to 2 minutes long each in HD resolution (720p), and corresponding dubbed audio in 2 or 3 different languages for each video, including French, Spanish, German and Italian.
Results. Table 3 shows the corresponding quantitative evaluation of the synthesized videos, lip synced to dubbed audio in different languages. Leveraging information from multiple identities, MI-NeRF slightly improves the LSE metrics and achieves similar visual quality with LipNeRF, at much lower training time (see Fig. 4). AD-NeRF (Guo et al., 2021) and GeneFace (Ye et al., 2023) lack in lip synchronization for this task. Fig. 5 compares the mouth position for two examples, lip synced to Spanish (original audio in English). The GAN-based method, Wav2Lip (Prajwal et al., 2020), frequently produces artifacts and blurry results. Since it operates in the 2D image space, it cannot handle large 3D movements. AD-NeRF overfits to the training audio and cannot generalize well to different speech inputs. LipNeRF performs well, but sometimes cannot produce unseen expressions (e.g. does not close the mouth in the first row), since it is only trained on a single video of limited duration. Both AD-NeRF and LipNeRF are standard single-identity NeRFs, requiring expensive identity-specific optimization.
@@ 4.5 Short-Video Personalization
MI-NeRF can also be adapted to an unseen identity, i.e. that is not seen as a part of the initial training set (see Sec. 3.4). Fig. 6 demonstrates the results when only a very short video of the new identity is available (1 or 3 seconds length). Please note that in this case we use a small number of consecutive frames, and thus a very small part of the expression space is covered. In this case, the single-identity INSTA (Zielonka et al., 2023) cannot learn a good head representation of the identity, given that very short video, and produces artifacts. HeadNeRF is trained on a large dataset of images and its fitting requires only a single image. However, it fails to capture the identity accurately. Our method demonstrates robustness under unseen expressions and novel views for the new identity. We include more quantitative and qualitative results for our short-video personalization in the appendix. We also encourage the readers to watch our supplementary video.
Figure 6: Short-Video Personalization. Learning a novel identity from a short video of 1 or 3 seconds duration. Qualitative comparison of MI-NeRF with the single-identity INSTA (Zielonka et al., 2023) and the NeRF-based parametric head model HeadNeRF (Hong et al., 2022). Our method demonstrates robustness across unseen expressions and novel views for any new identity.
@@ 5 Conclusion
In this work, we introduce MI-NeRF that learns a single dynamic NeRF from monocular talking face videos of multiple identities. We propose a multiplicative module that captures the non-linear interactions of identity and non-identity specific information. Trained on multiple identities, MI-NeRF significantly reduces the training time over standard single-identity NeRFs. Our model can be further personalized for a target identity, given only a short video of a few seconds length, achieving state-of-the-art performance for facial expression transfer and talking face video synthesis. In the future, we envision extending our approach to thousands of identities, learning collectively from very short in-the-wild video clips.
Reproducibility Statement
Our plan is to make the source code of our model publicly available once our work is accepted. We provide comprehensive documentation of the hyperparameters employed and offer detailed explanations of all the techniques used, supported by thorough ablation studies.
Appendix A Technical analysis of the models
@@
Notation: Tensors are symbolized by calligraphic letters, e.g., $\bm{\mathcal{X}}$. The mode-$m$ vector product of $\bm{\mathcal{X}}$ with a vector $\bm{u}\in\mathbb{R}^{I_{m}}$ is denoted by $\bm{\mathcal{X}}\times_{m}\bm{u}$. A core tool in our analysis is the CP decomposition (Kolda & Bader, 2009). By considering the mode-$1$ unfolding of an $M^{\text{th}}$-order tensor $\bm{\mathcal{X}}$, the CP decomposition can be written in matrix form as in Kolda & Bader (2009):
@@
A.1 Proof of Prop. 1
To prove the Prop. 1 (see Sec. 3.2 of the main paper), we will construct the special form from the general multiplicative interaction.
@@
where $\bm{W}_{2},\bm{W}_{3}\in\mathbb{R}^{o\times d}$ are learnable parameters. We can use the CP decomposition (Kolda & Bader, 2009) to induce a low-rank decomposition on the third-order tensor $\bm{\mathcal{W}}$. This results in the following expression:
@@
where * is the Hadamard product. If we set $o=k=d$ and set $A^{T}=\bm{U}_{1},B^{T}=\bm{U}_{2}$, then we end up with the model of Eq. 1 of the main paper, which concludes the proof.
A.2 Intuition on Eq. 2
The model of Eq. 2 can also capture multiplicative interactions in case of $N=2$ or high-degree interactions in case $N>2$. We provide some constructive proof to show the case of $N=2$ below, since the number of terms increases fast already for this case. Subsequently, we focus on the case of high-degree interactions.
Proposition 2.
For $N=2$, the model of Eq. 2 captures multiplicative interactions between the expression $\bm{e}$ and the identity $\bm{i}$ vector.
Proof.
For $N=2$, Eq. 2 becomes:
$\begin{split}H(\bm{e},\bm{i})=\bm{C}\left\{\left(\bm{U}_{(2,1)}\bm{e}+\bm{U}_{(2,2)}\bm{i}\right)*\left(\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}\right)\right\}+\bm{C}\left(\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}\right)=\\ \bm{C}\left\{(\bm{U}_{(2,1)}\bm{e})*(\bm{U}_{(1,1)}\bm{e})\right\}+\bm{C}\left\{(\bm{U}_{(2,1)}\bm{e})*(\bm{U}_{(1,2)}\bm{i})\right\}+\bm{C}\left\{(\bm{U}_{(2,2)}\bm{i})*(\bm{U}_{(1,1)}\bm{e})\right\}+\\ \bm{C}\left\{(\bm{U}_{(2,2)}\bm{i})*(\bm{U}_{(1,2)}\bm{i})\right\}+\bm{C}\bm{U}_{(1,1)}\bm{e}+\bm{C}\bm{U}_{(1,2)}\bm{i}\;.\end{split}$ (11)
Each of the first terms of the last equation arises from a CP decomposition with specific factor matrices (the inverse process from Eq. 8 to Eq. 9 can be followed). Therefore, Eq. 11 captures multiplicative interactions between the two variables, including second-degree interactions among the same variable. ∎
Proposition 3.
For $N>2$, the model of Eq. 2 captures high-degree interactions between the expression $\bm{e}$ and the identity $\bm{i}$ vector.
Proof.
The number of terms increase rapidly in Eq. 2. Without loss of generality, (a) we will only use the multiplicative term (and ignore the additive term of $+\bm{x}_{n-1}$, since this is not contributing to the higher-degree terms) and (b) we will showcase this for $N=3$. For $N=3$, we obtain the following formula:
@@
Following similar arguments as the proofs above and the unfolding of the CP decomposition, we can exhibit that all of those terms arise from CP decompositions with specific factor matrices, while each one captures a triplet of the form $(\bm{\tau}_{1},\bm{\tau}_{2},\bm{\tau}_{3})$ with $\bm{\tau}\in\{\bm{e},\bm{i}\}$. ∎
Appendix B Additional Ablation Study